# SPARTAN: A Sparse Transformer Learning Local Causation

### v2

@@ 6 Conclusion
We tackle the problem of adaptation in world models through the lens of local causal models. To this end, we propose SPARTAN, a structured world model that jointly performs dynamics model learning and causal discovery. We show on image-based datasets that attention mechanisms alone are not sufficient for discovering causal relationships and develop a novel sparsity regularisation scheme that learns accurate causal graphs. As a result, SPARTAN significantly extend the state-of-the-art with improved interpretability and few-shot adaptation capabilities while maintain the same level of prediction accuracy.
@@
There are several limitations to our approach, which serves as pointers for future investigations. 1) We empirically show that local causal graphs can be learnt from data. While we build on prior works in causal discovery that have theoretical guarantees [21], we do not make theoretical statements about the identifiability of local causal graphs in the setting where the scene composition can vary between each data sample. Future work should explore the conditions under which local graphs can be identified. 2) During adaptation, SPARTAN adapts by optimising over the learnt ’intervention space’. We have shown that this is sufficient to generalise to combinations of seen interventions or different number of objects. However, in more extreme cases where the test time environment contains completely new behaviours (e.g. ball teleporting), there might not be a corresponding intervention token in the space, while a finetuning approach would be able to converge to the right behaviour (given enough data). One avenue of exploration is to consider procedurally generating intervened environments, akin to domain randomisation [40, 54], to cover the space of all meaningful interventions. 3) Our approach relies on pre-disentangled object representations. An interesting extension of our work is to investigate whether local sparsity can induce the emergence of disentangled causal representations [33, 30] by jointly training an encoder with the dynamics model.
Acknowledgments and Disclosure of Funding
This research was supported by an EPSRC Programme Grant (EP/V000748/1). The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility (http://dx.doi.org/10.5281/zenodo.22558) and SCAN computing cluster in carrying out this work. Ingmar Posner holds concurrent appointments as a Professor of Applied AI at the University of Oxford and as an Amazon Scholar. This paper describes work performed at the University of Oxford and is not associated with Amazon.”
@@ A.1 Hyperparameters
SPARTAN and Transformer baseline are implemented as stacked transformer encoder layers with residual connections. The Global Graph baseline is implemented as an ensemble of MLPs with learnable adjacency matrix masks. This is similar [33, 20] with the modification that our baseline work on object embeddings rather than scalar states. Each MLP predicts a separate object embedding. For each MLP, the inputs object tokens are masked according to a learnable adjacency matrix before passing through the model. A sparse regularisation is applied to the adjacency matrix. The hyperparameters for SPARTAN and the baselines are shown in table 3 and 4. For the experiments on the simulated datasets, all models are trained on single GPUs (mixture of Nvidia V100 and RTX 6000) and converge within 3 days.
For the Traffic domain, predictions need to be multimodal since each vehicle has multiple possible trajectories. This is common for the dataset we consider. To overcome this, we use MTR [51] as the base architecture. MTR has two stages: first, it uses self-attention between tokenised map lines and vehicle trajectories; second, it applies cross attention using multiple queries, each representing one possible mode of motion. The output for each query is then used as a Gaussian mixture to predict multimodal motion patterns. We adapt SPARTAN to this architecture by swapping all attention layers with the sparsified version proposed in this paper. The original work also proposes a local attention masking scheme for each layer which only allow each token to attend to the k-nearest neighbour based on the position of each object. We refer to this as Local Attn in table 2. In the traffic domain, the models are trained in parallel on 4 GPUs due to the size of each scene (roughly 1000 tokens). Training takes less than one week for the baseline MTR model and under two weeks for SPARTAN.
@@
Optimiser Adam[26] Adam
@@ A.2 Lagrangian Relaxation
As discussed in Sec. 3.3, we formulate our training as a constrained optimisation problem. Specifically, we aim to minimise the expected number of edges and interventions while constraining the error to be lower than a target value,
@@
Applying Lagrangian relaxation, we obtain the following min-max objective,
@@
We solve this by taking gradient steps on $\theta$ and $\lambda$ in an alternating manner. To ensure $\lambda$ is positive, we perform updates in the form of $\lambda\leftarrow\alpha*exp(MSE-\tau)*\lambda$, where $\alpha$ is the step size. Intuitively, this has the effect of increasing $\lambda$ when the error is higher than the target, i.e. weighting the error term when the error is high. In practice, we initialise $\lambda$ to be high and set $\tau$ to be the error achieved by the fully connected model. We use a moving-average estimator for the $MSE-\tau$ term for stability. We also rewrite the training objective to $(MSE-\tau)+|\bar{A}|/\lambda$ so that the loss value is more stable. This acts as a scheduling scheme where the optimisation focuses on learning the dynamics and switch to pruning redundant edges once the error is below the target. Fig. 5 shows example training curves which demonstrate the training dynamics of SPARTAN. At the start of the training, $log(\lambda)$ stays low as the model prediction error improves. This allows the increase of the number of active edges. As the prediction error become low enough, the sparsity penalty automatically increases and the number of edges gradually decreases while maintaining the prediction accuracy.
@@ 4 Disabled cannon
The Interventional Pong environment is based on the Pong game and is originally developed in [36] for investigating causal representation learning. In the original work, the set of interventions included randomly perturbing the positions of the balls and the paddles. We modify the interventions to model dynamics changes, such as adding gravity or changing the way the ball bounces off the paddles, as these are more appropriate in the context of world model learning. A similar setup is also used in [20] in the context of few-shot adaptation. The original Interventional Pong dataset is under the BSD 3-Clause Clear License.
The environment contains 4 objects, the left paddle, the right paddle, the ball and the score. Here, local causal edges include ball $\rightarrow$ left, right paddles as the paddles follow the ball; paddles $\rightarrow$ ball when the ball collides with the paddles; ball $\rightarrow$ score when the ball crosses the left or right boundary lines. Observations are provided as 32 $\times$ 32 $\times$ 3 RGB images with masks for each object. To capture velocity information for ball, the ball image shows the position of the ball over 2 consecutive timesteps in different colours. Each object is separately encoded into 32-dimensional latent object-slots using a VAE [27]. The configurations of different intervened environments are shown in table 5.
@@ B.2 CREATE
[…]
@@ B.3 Traffic
We use the Waymo Open Dataset [52] for our Traffic domain. The dataset consists of real world scenes which include map lines (such as lane markings) and vehicle trajectories. The map lines are provided as polylines (i.e. each line is represented as a set of positions). The history of the state of each vehicle is provided as a list of positions, headings and velocities. Up to one second of state is given for each vehicle and the task is to predict 9 seconds into the future for the centre vehicle. The number of map lines and vehicle trajectories are different for each scene, with roughly 800 map lines and 50 vehicles per scene. This presents a major challenge for prior approaches in causal world model learning. Since the dataset is not generated by simulation, we do not apply interventions to the data.

### v3

@@ 6 Conclusion
We tackle the problem of adaptation in world models through the lens of local causal models. To this end, we propose SPARTAN, a structured world model that jointly performs dynamics model learning and causal discovery. We show on image-based datasets that attention mechanisms alone are not sufficient for discovering causal relationships and develop a novel sparsity regularisation scheme that learns accurate causal graphs, resulting in significantly improved interpretability and few-shot adaptation capabilities without compromising prediction accuracy.
@@
There are several limitations to our approach that serve as pointers for future investigations. 1) We empirically show that local causal graphs can be learnt from data. While the use of sparsity regularisation is grounded in prior works that have theoretical guarantees [hwang2024finegrained], we do not make theoretical statements about the identifiability of local causal graphs in the setting where the scene composition can vary between each data sample. Future work should explore the conditions under which local graphs can be identified. 2) During adaptation, SPARTAN adapts by optimising over the learnt ’intervention space’. We have shown that this is sufficient to generalise to combinations of seen interventions or different numbers of objects. However, in more extreme cases where the test time environment contains completely new behaviours (e.g. ball teleporting), there might not be a corresponding intervention token in the space, while a finetuning approach would be able to converge to the right behaviour (given enough data). One avenue of exploration is to consider procedurally generating intervened environments, akin to domain randomisation [peng2018sim, tobin2017domain], to cover the space of all meaningful interventions. 3) Our approach relies on pre-disentangled object representations. Note that while we have used ground-truth segmentation masks for objects in our evaluations, pre-trained object-centric representations can also be utilised. We provide further discussion on this in App. D. An interesting extension of our work is to investigate whether local sparsity can induce the emergence of disentangled causal representations [lei2023variational, lachapelle2022disentanglement] by jointly training an encoder with the dynamics model.
Acknowledgments
This research was supported by an EPSRC Programme Grant (EP/V000748/1). We would also like to thank the University of Oxford Advanced Research Computing (ARC) (http://dx.doi.org/10.5281/zenodo.22558) and the SCAN facility in carrying out this work.
@@ A.1 Hyperparameters
SPARTAN and the Transformer baseline are implemented as stacked transformer encoder layers with residual connections. The Global Graph baseline is implemented as an ensemble of MLPs with learnable adjacency matrix masks. This is similar [lei2023variational, huang2021adarl] to the modification that our baseline works on object embeddings rather than scalar states. Each MLP predicts a separate object embedding. For each MLP, the input object tokens are masked according to a learnable adjacency matrix before being passed through the model. A sparse regularisation is applied to the adjacency matrix. The hyperparameters for SPARTAN and the baselines are shown in table 3 and 4. For the experiments on the simulated datasets, all models are trained on single GPUs (mixture of Nvidia V100 and RTX 6000) and converge within 3 days.
For the Traffic domain, predictions need to be multimodal since each vehicle has multiple possible trajectories. This is common for the dataset we consider. To overcome this, we use MTR [shi2022motion] as the base architecture. MTR has two stages: first, it uses self-attention between tokenised map lines and vehicle trajectories; second, it applies cross attention using multiple queries, each representing one possible mode of motion. The output for each query is then used as a Gaussian mixture model to predict multimodal motion patterns. We adapt SPARTAN to this architecture by swapping all attention layers with the sparsified version proposed in this paper. The original work also proposes a local attention masking scheme for each layer, which only allows each token to attend to the k-nearest neighbour based on the position of each object. We refer to this as Local Attn in table 2. In the traffic domain, the models are trained in parallel on 4 GPUs due to the size of each scene (roughly 1000 tokens). Training takes less than one week for the baseline MTR model and under two weeks for SPARTAN.
@@
Optimiser Adam[adam] Adam
@@ A.2 Lagrangian Relaxation
As discussed in Sec. 3.3, the model is sensitive to the choice of $\lambda$. Since removing non-causal edges should not deteriorate the prediction accuracy, we formulate a constrained optimisation problem where we minimise $|\bar{A}|$ under the constraint that $\mathcal{L}\leq\tau$, where $\tau$ is the target loss,
@@
We set the target as the loss achieved by a fully connected model. Applying Lagrangian relaxation [bertsekas2014constrained], we obtain the following min-max objective,
@@
We solve this by taking gradient steps on $\theta$ and $\lambda$ in an alternating manner. To ensure $\lambda$ is positive, we perform updates in the form of $\lambda\leftarrow\alpha*exp(MSE-\tau)*\lambda$, where $\alpha$ is the step size. Intuitively, this has the effect of increasing $\lambda$ when the error is higher than the target, i.e. weighting the error term when the error is high. This is analogous to the GECO [rezende2018taming] framework for tuning the KL loss in a VAE. In practice, we initialise $\lambda$ to be high and set $\tau$ to be the error achieved by the fully connected model. We use a moving-average estimator for the $MSE-\tau$ term for stability. We also rewrite the training objective to $(MSE-\tau)+|\bar{A}|/\lambda$ so that the loss value is more stable. This acts as a scheduling scheme where the optimisation focusses on learning the dynamics and switches to pruning redundant edges once the error is below the target. Fig. 5 shows example training curves demonstrating the training dynamics of SPARTAN. At the start of the training, $log(\lambda)$ remains low as the model prediction error improves. This allows an increase in the number of active edges. As the prediction error becomes low enough, the sparsity penalty automatically increases, and the number of edges gradually decreases while maintaining prediction accuracy.
@@ 4 Disabled cannon
The Interventional Pong environment is based on the Pong game and was originally developed in [lippe2022citris] for investigating causal representation learning. In the original work, the set of interventions included randomly perturbing the positions of the balls and the paddles. We modify the interventions to model dynamics changes, such as adding gravity or changing the way the ball bounces off the paddles, as these are more appropriate in the context of world model learning. A similar setup is also used in [huang2021adarl] in the context of few-shot adaptation. The original Interventional Pong dataset is under the BSD 3-Clause Clear License.
The environment contains 4 objects: the left paddle, the right paddle, the ball, and the score. Here, local causal edges include ball $\rightarrow$ left, right paddles as the paddles follow the ball; paddles $\rightarrow$ ball when the ball collides with the paddles; ball $\rightarrow$ score when the ball crosses the left or right boundary lines. Observations are provided as 32 $\times$ 32 $\times$ 3 RGB images with masks for each object. To capture velocity information for the ball, the ball image shows the position of the ball over 2 consecutive timesteps in different colours. Each object is separately encoded into 32-dimensional latent object-slots using a VAE [kingma2022autoencoding]. The configurations of different intervened environments are shown in table 5.
@@ B.2 CREATE
The CREATE environment is built using the CREATE simulator [jain2020general] which consists of physical interactions between various objects. The simulator is under the MIT License. In our experiments, the objects in the environment include {blue, red, green} balls, a wall, a ladder, and a cannon. The balls collide elastically. All objects are initialised at random positions with random orientations. The wall remains static for each sequence. The ladder object allows the balls to climb up. The cannon object projects the balls at a fixed velocity in the direction it points at. Here, local edges capture when the objects collide or interact with each other. Observations are given as 84 $\times$ 84 $\times$ 3 images with masks for each object. To capture the velocities of the balls, a trail of the past positions (up to two timesteps) is shown in the image for each ball. The intervened environments are shown in table 6. In the unseen environments, we test the models using different scene composition, i.e. different numbers of objects.
@@ B.3 Traffic
We use the Waymo Open Dataset [waymo] for our Traffic domain. The dataset consists of real world scenes, which include map lines (such as lane markings) and vehicle trajectories. The map lines are provided as polylines (i.e. each line is represented as a set of positions). The history of the state of each vehicle is provided as a list of positions, headings and velocities. Up to one second of state is given for each vehicle, and the task is to predict 9 seconds into the future for the centre vehicle. The number of map lines and vehicle trajectories is different for each scene, with roughly 800 map lines and 50 vehicles per scene. This presents a major challenge for prior approaches in causal world model learning. Since the dataset is not generated by simulation, we do not apply interventions to the data.
Appendix C Relation to existing approaches and extra baselines
The aim of this work is to integrate the inductive bias of sparse interactions, common in causal discovery, into a transformer-based architecture. As such, we focus our evaluations on environments that are commensurate in complexity and scale to the settings on which prior transformer world models are evaluated. While there exist prior approaches for local causal learning, as discussed in the related works section, these methods cannot be readily scaled to the environments in the experiments presented in this paper. This is due to two principal reasons: first, prior methods operate directly on state-based observations, such as x, y positions of objects, whereas in our environments, objects are represented as learnt embeddings (typically 64 dimensions) that are mapped from images; second, and more fundamentally, our environments, in particular CREATE and traffic domains, require the world model to operate on different numbers of objects and different compositions of objects across samples, e.g. different numbers of vehicles across scenes. A transformer-based world model can accommodate these requirements by treating each object embedding as a separate token. Here, we discuss the particular architectural choices of prior works that hinder their application in these environments.
ELDEN [ELDEN] explores the role of local causal models in the context of intrinsic motivations for RL. Here, the forward dynamics model is trained with a sparsity regularisaion, proxied by using the L1 norm, on the partial derivatives between the states. In principle, this technique can be extended beyond scalar states to the case in which each object is represented by higher dimensional embedding vectors. However, we implemented ELDEN for our experiments and found that, since the loss is a function of many partial derivatives, the optimisation process becomes prohibitively memory-intensive and highly unstable.
FCDL [hwang2024finegrained] investigates the robustness improvements afforded by local causal models. This approach learns a codebook of possible local causal graphs and infers the correct structure conditioned on the state. Sparse regularisation is then applied to the set of local causal graphs represented by the codebook. The drawback of this approach is that the local graphs must have the same set of nodes across samples, meaning that all scenes must have exactly the same composition of objects.
CAI [seitzer2021causal] considers a related but different setting from our method, where the goal is to estimate the causal influence of an agent on the states of the objects in the scene. This approach relies on testing the conditional independence using conditional mutual information estimation between the agent’s action and the objects based on a learned dynamics model. However, estimating conditional mutual information is challenging in the multivariate case.
CODA [pitis2020count] uses threshold heuristics to determine local structures. We compare against this in our main experiments and show that attention alone is not able to reliably learn the correct dependency structure.
Table 7: Average SHD between the learned graphs and the ground-truth. Lower is better.
SPARTAN ACD Sparse ACD
Intervention Pong 1.50 10.27 2.67
Create 1.17 39.77 5.45
Beyond the local causal structure learning literature, the idea of learning context-dependent causal structures is also explored in Amortised Causal Discovery (ACD) [lowe2022amortized] which uses a GNN backbone to infer causal structures conditioned on the history of states in a time-series setting. This approach can be adapted to the state-dependent local causal structure case by conditioning the graph encoder on the current state rather than on an entire time-series. In table 7, we compare the graph learning accuracy between SPARTAN and ACD. We note that ACD requires setting a prior level of expected sparsity. Here, we compare against a naive choice of 50% sparsity (labelled ACD) as well as the sparse baseline, which uses the ground-truth sparsity as a prior (labelled Sparse ACD). The results show that SPARTAN significantly outperforms this baseline, even if the true sparsity level, which is in practice not accessible, is specified. This further corroborates the efficacy of our model.
Appendix D Using pretrained object-centric representations
Many existing works that aim to capture object interactions, including SlotFormer [wu2023slotformer], use pre-trained object-centric representations [slot_attention] that handle the segmentation and tracking problems. Our approach can be readily applied to any such pretrained representations. In our experiments, we opted to use ground-truth masked object representations for the principal reason that they enable the evaluation of the learned graphs: quantitatively evaluating the SHD of learned graphs requires mapping each latent embedding to the corresponding ground-truth object. Learned object embeddings can make this mapping noisy and ambiguous due to imperfect segmentations. Since learning high-quality representations is not the main focus of this work, we use ground-truth masked object representations to ensure the fairness and validity of our evaluation. To demonstrate the generality of our results, we have trained our model using a learned slot representations in the pong environment. Here, to measure the SHD between the learned graph and the ground-truth causal graph, we use the visualised attention masks from Slot Attention to map the latent object representations to ground-truth objects. In table 8, we present the results, which are consistent with the main findings in the paper: our model can achieve the same level of prediction while capturing the underlying graph structure more accurately.
Table 8: The prediction error and SHD of the learned graphs in the Intervention Pong environment.
SPARTAN Transformer
Prediction Error ($\pm$ SE) 7.21 $\pm$ 0.69 6.90 $\pm$ 0.83
SHD ($\pm$ SE) 2.69 $\pm$ 0.02 8.52 $\pm$ 0.02